# Can Carbon-Aware Electric Load Shifting Reduce Emissions? An Equilibrium-Based Analysis

### v1

@@
Abstract
An increasing number of electric loads, such as hydrogen producers or data centers, can be characterized as carbon-sensitive, meaning that they are willing to adapt the timing and/or location of their electricity usage in order to minimize carbon footprints. However, the emission reduction efforts of these carbon-sensitive loads rely on carbon intensity information such as average carbon emissions, and it is unclear whether load shifting based on these signals effectively reduces carbon emissions. To address this open question, we investigate the impact of carbon-sensitive consumers using equilibrium analysis. Specifically, we expand the commonly used equilibrium model for electricity market clearing to include carbon-sensitive consumers that adapt their consumption based on an average carbon intensity signal. This analysis represents an idealized situation for carbon-sensitive loads, where their carbon preferences are reflected directly in the market clearing, and contrasts with current practice where carbon intensity signals only become known to consumers aposteriori (i.e. after the market has already been cleared). We include both illustrative examples and larger numerical simulations, including benchmarking with other methods, to illuminate the contributions and limitations of carbon-sensitive loads in power system emission reductions.
Index Terms:
@@ Abstract
I Introduction
@@
The Paris Agreement recognizes emissions abatement as a critical strategy for mitigating global warming. Due to both the significant emissions from electricity generation and the potential for reducing these emissions by increasing renewable and low-carbon generation technologies, the electricity sector has been at the forefront of implementing decarbonization policies. However, the inherent variability of renewable generation, which is highly dependent on weather conditions (e.g., wind speed, solar radiation), necessitates aligning electricity demand with supply availability to maintain reliability. Demand-side flexibility is a crucial component of this alignment, i.e. consumers must be incentivized to adjust their consumption in response to renewable generation surpluses or deficits. At the same time, a growing number of carbon-sensitive loads, such as data centers and hydrogen production facilities, are emerging. These entities actively optimize the timing and location of their electricity usage to reduce carbon footprints, driven either by pledges to environmental responsibility or economic incentives such as production tax credits. However, many of these consumers are unwilling to give up control of their power consumption e.g. by participating in a market, and instead prefer to locally optimize their consumption based on externally provided information such as carbon intensity signals. In this paper, we try to explore the coordination of carbon-sensitive loads with the rest of the electricity market clearing through a novel equilibrium formulation for an electricity market clearing where we model the impact of average carbon emissions on consumers’ decision making.
@@
Whereas calculating and potentially penalizing carbon emissions from generation is a relatively straightforward task, the question of how to allocate emissions from a group of generators across a group of consumers is less obvious and subject to ongoing debate [1]. A common approach is to define a carbon intensity signal (with the most commonly used and widely available signal being average carbon emissions [1]) which is computed based on the average rate of emissions per megawatt-hour (MWh) of electricity consumed [2, 3]. To differentiate the impacts of locations on the carbon reduction, some researchers have advocated for the use of locational marginal carbon emissions [4, 5, 6, 7, 8], which provides information about how a change in the electricity consumption at a given location would impact total emissions. Other, more recently proposed metrics include locational average carbon emissions [9], which is based on the concept of carbon flow [10, 11], and adjusted locational marginal emissions [8], which combines characteristics of locational and average emissions.
@@
While carbon intensity signals are generally used for carbon accounting, e.g. as outlined in [1], carbon-sensitive loads also leverage these signals to decide where and/or when to consume electricity. In practice, such adjustment would happen in a sequential manner, where the electricity market is first cleared assuming a given load profile and carbon intensity signals are then calculated based on the market outcome. Subsequently, carbon-sensitive consumers utilize these carbon intensity signals to adapt their consumption across time and space, with the goal of achieving carbon reductions. The impact of these changes have been studied in the context of real-time market clearing [7, 12], where the updated load profile becomes the input to the next market clearing, and in the context of day-ahead market clearing [8], where the updated load profile is taken into account in the intra-day market. As observed in [7, 12, 8], this sequential approach leads to suboptimal outcomes because consumers act on carbon intensity signals arising from the previous market clearing, without insight into how their (and other consumers/generators) actions may change the generation dispatch and potentially impact the carbon intensity signals in the next market clearing.
@@
Our recent work in [13] proposes a novel electricity market clearing model that incorporates the allocation of emissions from generations to loads and allows for consideration of consumer-side carbon costs. Continuing to further this line of research, in this paper, we propose a game-theoretic model, where all market participants jointly identify a solution to the market clearing by solving an equilibrium problem that includes average carbon signal and their impacts on electricity consumption. As we mentioned before, there are several different methods to define the carbon emission intensity of electricity consumption. We choose to use average carbon signals because this metric is widely available, both from system operators and companies [14, 15], and because it is commonly used for carbon accounting under the Greenhouse Gas Protocol [16, 1].
@@
The major contributions of this paper are twofold. First, we propose a novel equilibrium model considering the average carbon signal, which assumes that consumers have “perfect” knowledge of their own carbon emission intensity. This formulation provides a new perspective to evaluate the effectiveness of different carbon signals and contributes to new carbon signal design. To the best of our knowledge, this is the first paper that investigates the impact of carbon-sensitive consumers on carbon reductions from a game theoretic perspective. Second, we demonstrate how carbon-sensitive consumers behave in the equilibrium model against the sequential model. Based on the real-world IEEE RTS-GMLC system, we then compare the equilibrium model with benchmark models with or without considering the impact of carbon emissions.
@@
The remainder of the paper is organized as follows: Section II describes our proposed equilibrium model. In Section III, we show results and conclusions from our numerical study, while Section IV summarizes and concludes.
II Carbon-aware Equilibrium Electricity Market Clearing Model
@@
In this section, we introduce the proposed equilibrium problem formulation. After that, we employ a simplified three-bus system to illustrate how loads adapting to carbon signals impact the generation dispatch and carbon emissions of the overall power system.
@@
However, before diving into the details of our formulation, we first introduce some basic background on equilibrium models in the electric sector. Equilibrium models, or game-theoretic models, are commonly employed to investigate strategic behavior in deregulated electricity markets [17, 18]. In such models, there is typically three (in the electricity pool model, which ignores transmission constraints) or four (in the transmission-constrained model) types of participants, namely generators, consumers, transmission owners, and ISO. In the traditional carbon-agnostic electricity market, generators, consumers, and transmission owners aim at maximizing their profits given electricity prices as input parameters, and the ISO determines the market clearing price given power generation and electricity consumption as input parameters. All participants are, between them, playing a noncooperative game, and thus the optimal solution to this game is defined as a Nash equilibrium, corresponding to a situation where no participant can improve their outcomes by unilaterally changing their decisions. Under a price-taking assumption (no strategic bidding or market power) the Nash equilibrium problem is equivalent to a standard market clearing optimization problem, i.e., solutions from the equilibrium problem are the same as those from the optimization problem. The equilibrium can often be rewritten as a complementarity problem; more details, examples and formulations can be found in [19, 20, 21].
II-A Model Formulation
@@
We next introduce our proposed equilibrium formulation with carbon-sensitive loads. As in the typical equilibrium model for electricity markets, we consider generators, consumers, transmission owners, and the ISO as participants. As is common, generators determine their optimal power generation based on electricity prices and the transmission owner maximizes their profit by buying power at a bus and selling it back at another one. Different from the standard model, consumers determine their electricity consumption considering both electricity prices and carbon emissions, and the ISO determines both the price and the average carbon signal. Below, we provide the parametric optimization problems solved by the generators, consumers, and transmission owners, respectively, and the complementarity conditions involving the ISO.
@@
We consider an electric power network with the set of nodes, loads, transmission lines, and generators denoted $\mathcal{N}$, $\mathcal{D}$, $\mathcal{L}$ and $\mathcal{G}$, respectively. Let $\mathcal{G}_{i}\subset\mathcal{G}$ and $\mathcal{D}_{i}\subset\mathcal{D}$ be the subset of generators and loads connected to the node $i$, and $(i,j)\in\mathcal{L}$ denote the transmission line from node $i\in\mathcal{N}$ to node $j\in\mathcal{N}$. Let $g\in\mathcal{G}$ and $d\in\mathcal{D}$ denote the indexes for generators and consumers, respectively.
@@
Consumers: Each consumer in our model maximizes their benefits of consuming electricity, but different from a standard formulation, the consumers consider both price and carbon emissions in their decisions. This gives rise to a carbon-dependent consumer utility, which we define as $r_{D,d}-\lambda\cdot c_{D,d}$. Here, $r_{D,d}$ is the revenue for consumers derived from consuming electricity, $c_{D,d}$ denotes the carbon cost, given in units of [\$/tons $CO_{2}$] and $\lambda$ is the value of the average carbon signal. Specifically, given price $p_{i}$ and the average carbon signal $\lambda$ calculated by ISO in the later part, each consumer solves the following problem:
$\max_{P_{D,d}}$ $(r_{D,d}-p_{i:d\in\mathcal{D}_{i}}-\lambda\cdot c_{D,d})\cdot P_{D,d}$ (2a)
$s.t.$ $P_{D,d}^{\min}\leq P_{D,d}\leq P_{D,d}^{\max},$ (2b)
@@
where $p_{i:d\in\mathcal{D}_{i}}$ represents that the electricity price at the node $i$ to which the consumer $d$ is connected. Each consumer incorporates the carbon term $\lambda\cdot c_{D,d}$ into their utility evaluation. In practice, the cost parameter $c_{D,d}$ may be directly tied to concrete costs, such as carbon emission penalties from carbon taxes or cap-and-trade schemes, or could be an internally set “carbon cost”, reflecting how much revenue the consumer is willing to forgo to avoid carbon emissions. Note that besides using carbon-dependent utilities to quantity the impacts of carbon emissions on consumers’ decision-making, there are also other ways to quantify this impact, e.g., setting carbon emission limits [22, 23]. However, it can be very challenging for consumers to define what the values of the carbon emission cap should be.
@@
Transmission Owners: Transmission owners act as spatial arbitragers and aim to maximize the power transmission profit given electricity prices $p_{i}$. Specifically, the transmission owner maximizes their profit by buying power at a bus and selling it back at another one (arbitrage through electricity price difference between two bus nodes), which is characterized by the following mathematical optimization problem:
$\max_{\theta_{i}}$ $\sum_{i}p_{i}\cdot\left(\sum_{j,(i,j)\in\mathcal{L}}\beta_{ij}(\theta_{j}-\theta_{i})\right)$ (3a)
$s.t.$ $-F_{ij}^{\rm{lim}}\leq\beta_{ij}(\theta_{i}-\theta_{j})\leq F_{ij}^{\rm{lim}},\quad\forall(i,j)\in\mathcal{L},$ (3b)
$\theta_{ref}=0,$ (3c)
@@
where $\beta_{ij}\in\mathbb{R}$ denoting the susceptance value of the transmission line $(i,j)$ from node $i$ to node $j$. Constraints (3b) are the transmission line limits, where $F_{ij}^{\rm{lim}}$ represents the transmission capacity (which we assume is the same in both directions). Constraint (3c) sets the voltage angle at the reference node to zero.
@@
ISO: The ISO clears the market by determining the electricity trading price at each node and provides the average carbon signal. Specifically, they set each nodal electricity price $p_{i}$ such that the supply-demand power balance
$\sum_{d\in\mathcal{D}_{i}}P_{D,d}+\sum_{j:(i,j)\in\mathcal{L}}\beta_{ij}(\theta_{i}-\theta_{j})=\sum_{g\in\mathcal{G}_{i}}P_{G,g}$ (4)
@@
$\lambda\sum_{d\in\mathcal{D}}P_{D,d}=\sum_{g\in\mathcal{G}}e_{G,g}\cdot P_{G,g},$ (5)
@@
where $e_{G}=\{e_{G,g}:\forall g\in\mathcal{G}\}$ is the carbon emission intensity vector for all generators (for generators using renewable sources, the carbon emission intensity values are zero).
@@
Note that (4) and (5) are written here as equality constraints but form a square complementarity system when paired with $p$ and $\lambda$, respectively. The number of equations in (4) matches the size of the variable $p$; similarly with (5) and $\lambda$. Both $p$ and $\lambda$ are parameters in the optimization problems (1)–(3).
The overall Nash equilibrium problem consists of the collection of parametric optimization problems (1), (2), (3) and the parametric complementarity conditions (4) and (5).
@@
Remark 2: When all $c_{D}=0$, our equilibrium problem reduces to the standard social welfare maximization problem. The equilibrium framework seeks to coordinate all participants toward a consistent set of values for $P_{G}$, $P_{D}$, $p$, and $\lambda$ that simultaneously ensures power balance and aligns with the optimal solutions of each participant’s individual optimization problem. Unlike the sequential method, where inaccuracies may arise from relying on lagged market-clearing signals (see [8, 7, 12]), the equilibrium method uses synchronized decision-making based on perfect real-time carbon signals, which we believe is better for analyzing the effectiveness of different carbon signals. We note that, similar to the standard market-clearing optimization problem, the equilibrium problem may have multiple solutions (e.g., when multiple generators at a node share identical costs).
II-B Three-bus Example Illustration

### v2

@@
Abstract.
An increasing number of electric loads, such as hydrogen producers or data centers, can be characterized as carbon-sensitive, meaning that they are willing to adapt the timing and/or location of their electricity usage in order to minimize carbon footprints. However, the emission reduction efforts of these carbon-sensitive loads rely on carbon intensity information such as average carbon emissions, and it is unclear whether load shifting based on these signals effectively reduces carbon emissions. To address this open question, we design a carbon-aware equilibrium model, which expands the commonly used equilibrium model for standard (carbon-agnostic) electricity market clearing to include carbon-sensitive consumers that adapt their consumption based on average carbon emission signals and carbon costs. This analysis represents an idealized situation for carbon-sensitive consumers, where their carbon preferences are reflected directly in the market clearing, and contrasts with current practice, where carbon emission signals only become known to consumers a posteriori (i.e., after the market has already been cleared). Furthermore, we extend our model to consider temporal load shifting and time-varying maximum renewable generations. We employ illustrative three-bus examples and numerical simulations on the IEEE RTS-GMLC system to reveal the limitations of the widely adopted average carbon emission signal for guiding carbon emission reduction. Our model offers a novel perspective for evaluating the effectiveness of different carbon signals and contributes to new carbon signal design.
@@ Abstract.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: The ACM International Conference on Future and Sustainable Energy Systems; June 23–-25, 2026; Banff, Canada isbn: 978-1-4503-XXXX-X/2026/06ccs: Hardwareccs: Hardware Power and energyccs: Hardware Renewable energyccs: Hardware Impact on the environment
1. Introduction
@@
The Paris Agreement recognizes emissions abatement as a critical strategy for mitigating global warming. Due to both the significant emissions from electricity generation and the potential for reducing these emissions by increasing renewable and low-carbon generation technologies, the electricity sector has been at the forefront of implementing decarbonization policies. However, the inherent variability of renewable generation, which is highly dependent on weather conditions (e.g., wind speed, solar radiation), necessitates aligning electricity demand with supply availability to maintain reliability. Demand-side flexibility is a crucial component of this alignment, i.e., consumers must be incentivized to adjust their consumption in response to renewable generation surpluses or deficits. At the same time, a growing number of carbon-sensitive loads, such as data centers and hydrogen production facilities, are emerging. These entities actively optimize the timing and location of their electricity usage to reduce carbon footprints, driven either by pledges to environmental responsibility or economic incentives such as production tax credits. However, many of these consumers are unwilling to give up control of their power consumption, e.g., by participating in a market, and instead prefer to locally optimize their consumption based on externally provided information such as carbon intensity signals. In this paper, we try to explore the coordination of carbon-sensitive loads with the rest of the electricity market clearing through a novel equilibrium formulation for an electricity market clearing where we model the impact of average carbon emissions on consumers’ decision-making.
1.1. Related Works
To mitigate carbon emissions in power systems, researchers and policy- makers have considered cap-and-trade emissions trading schemes (European Union, 2025; California Air Resource Board, 2025; International Carbon Action Partnership, 2025) and carbon tax schemes (Murray and Rivers, 2015; World Bank Group., 2025). Emissions trading provides greater environmental certainty in controlling overall emissions compared to a carbon tax, which defines a fixed emission price without restricting the quantity of GHGs emitted over a certain period (International Carbon Action Partnership, 2025; Stavins, 2019; Goulder and Schein, 2013). Many differences between these two schemes fade with specific implementation decisions (Stavins, 2019). Generally, these mechanisms reduce system-wide emissions in power systems by mainly increasing the cost of carbon-intense generation, thus dispatching more low-carbon generation. However, the increased generation costs are ultimately paid by consumers via elevated electricity prices (Nazifi et al., 2021). An analysis of the European emissions trading system showed that emission reductions primarily arise from reductions in demand due to higher electricity prices, rather than increased dispatch of clean generation (Chen et al., 2008; Sijm et al., 2005). Further, it is worth noting that a uniform carbon tax or the cap-and-trade emission trading scheme cannot account for differences in individual consumers’ preference for low-carbon power.
@@ 1.1. Related Works
Whereas calculating and potentially penalizing carbon emissions from generation is a relatively straightforward task, the question of how to allocate emissions from a group of generators across a group of consumers is less obvious (Bushnell, 2008) and subject to ongoing debate (Standard, 2014). A common approach is to define a carbon intensity signal (with the most commonly used and widely available signal being average carbon emissions (Standard, 2014)) which is computed based on the average rate of emissions per megawatt-hour (MWh) of electricity consumed (Bettle et al., 2006; Maps, 2025). To differentiate the impacts of locations on the carbon reduction, some researchers have advocated for the use of locational marginal carbon emissions (Hawkes, 2014; Siler-Evans et al., 2012; Goldsworthy and Aryai, 2023; Lindberg et al., 2022; Gorka et al., 2025), which provides information about how a change in the electricity consumption at a given location would impact total emissions. Other, more recently developed metrics include locational average carbon emissions (Chen et al., 2024a), which is based on the concept of carbon flow (Kang et al., 2015; Chen et al., 2024b), and adjusted locational marginal emissions (Gorka et al., 2025), which combines characteristics of locational and average emissions.
@@
While carbon intensity signals are generally used for carbon accounting, e.g., as outlined in (Standard, 2014), carbon-sensitive loads also leverage these signals to decide where and/or when to consume electricity. In practice, such adjustment would happen in a sequential manner, where the electricity market is first cleared assuming a given load profile and carbon intensity signals are then calculated based on the market outcome. Subsequently, carbon-sensitive consumers utilize these carbon intensity signals to adapt their consumption across time and space, with the goal of achieving carbon reductions. The impact of these changes has been studied in the context of real-time market clearing (Lindberg et al., 2022, 2021a), where the updated load profile becomes the input to the next market clearing, and in the context of day-ahead market clearing (Gorka et al., 2025), where the updated load profile is taken into account in the intra-day market. As observed in (Lindberg et al., 2022, 2021a; Gorka et al., 2025), this sequential approach leads to suboptimal outcomes because consumers act on carbon intensity signals arising from the previous market clearing, without insight into how their (and other consumers/generators) actions may change the generation dispatch and potentially impact the carbon intensity signals in the next market clearing.
@@
Our recent work in (Jiang and Roald, 2025) develops a novel electricity market clearing model that incorporates the allocation of emissions from generations to loads and allows for consideration of consumer-side carbon costs. Continuing to further this line of research, in this paper, we design a game-theoretic model where all market participants jointly identify a solution to the market clearing by solving an equilibrium problem that includes average carbon signal and their impacts on electricity consumption. As we mentioned before, there are several different methods to define the carbon emission intensity of electricity consumption. We choose to use average carbon signals because this metric is widely available, both from system operators and companies (CAISO, 2024; Tomorrow, 2024), and because it is commonly used for carbon accounting under the Greenhouse Gas Protocol (WBCSD and WRI, 2024; Standard, 2014).
1.2. Contributions
The major contributions of this paper are threefold.
@@ 1.2. Contributions
  • We develop a novel equilibrium model considering the average carbon signal, which assumes that consumers have “perfect” knowledge of their own carbon emission intensity. This formulation provides a new perspective to evaluate the effectiveness of different carbon signals and contributes to new carbon signal design. To the best of our knowledge, this is the first paper that investigates the impact of carbon-sensitive consumers on carbon reductions from a game-theoretic perspective.
  • We extend our equilibrium model to consider time-varying maximum renewable power availability and temporal load shifting, which enables the analysis of how average carbon signals impact temporal load shifting results when the total demand for each consumer is fixed across all time periods.
  • We demonstrate how carbon-sensitive consumers behave in the equilibrium model against benchmark models based on the IEEE RTS-GMLC system. Numerical results show that the average carbon intensity signal may not be the “right” signal for load shifting to achieve effective carbon reductions.
@@
The remainder of the paper is organized as follows: Section 2 describes our equilibrium model. Section 3 extends our model by considering time-varying maximum renewable generation and temporal load shifting. In Section 4, we show results and conclusions from our numerical study, while Section 5 summarizes and concludes.
2. Carbon-aware Equilibrium Electricity Market Clearing Model
@@
In this section, we introduce the equilibrium problem formulation. After that, we employ a simplified three-bus system to illustrate how loads adapting to carbon signals impact the generation dispatch and carbon emissions of the overall power system.
@@
However, before diving into the details of our formulation, we first introduce some basic background on equilibrium models in the electric sector. Equilibrium models, or game-theoretic models, are commonly employed to investigate strategic behavior in deregulated electricity markets (Ventosa et al., 2005; Hobbs and Pang, 2007). In such models, there are typically three (in the electricity pool model, which ignores transmission constraints) or four (in the transmission-constrained model) types of participants, namely generators, consumers, transmission owners, and ISO. In the traditional carbon-agnostic electricity market, generators, consumers, and transmission owners aim at maximizing their profits given electricity prices as input parameters, and the ISO determines the market clearing price given power generation and electricity consumption as input parameters. All participants are, between them, playing a noncooperative game, and thus the optimal solution to this game is defined as a Nash equilibrium, corresponding to a situation where no participant can improve their outcomes by unilaterally changing their decisions. Under a price-taking assumption (no strategic bidding or market power), the Nash equilibrium problem is equivalent to a standard market-clearing optimization problem, i.e., solutions from the equilibrium problem are the same as those from the optimization problem. The equilibrium can often be rewritten as a complementarity problem; more details, examples and formulations can be found in (Ferris and Pang, 1997; Gabriel et al., 2012; Ferris and Philpott, 2025).
2.1. Model Formulation
@@
We next introduce our equilibrium formulation with carbon-sensitive loads. As in the typical equilibrium model for electricity markets, we consider generators, consumers, transmission owners, and the ISO as participants. As is common, generators determine their optimal power generation based on electricity prices and the transmission owner maximizes their profit by buying power at one bus and selling it back at another one. Different from the standard model, consumers determine their electricity consumption considering both electricity prices and carbon emissions, and the ISO determines both the price and the average carbon signal. Below, we provide the parametric optimization problems solved by the generators, consumers, and transmission owners, respectively, and the complementarity conditions involving the ISO.
@@
We consider an electric power network with the set of nodes, loads, transmission lines, and generators denoted by $\mathcal{N}$, $\mathcal{D}$, $\mathcal{L}$ and $\mathcal{G}$, respectively. Let $\mathcal{G}_{i}\subset\mathcal{G}$ and $\mathcal{D}_{i}\subset\mathcal{D}$ be the subset of generators and loads connected to the node $i$, and $(i,j)\in\mathcal{L}$ denote the transmission line from node $i\in\mathcal{N}$ to node $j\in\mathcal{N}$. Let $g\in\mathcal{G}$ and $d\in\mathcal{D}$ denote the indexes for generators and consumers, respectively.
@@
Consumers: Each consumer in our model maximizes their benefits of consuming electricity, but different from a standard formulation, the consumers consider both price and carbon emissions in their decisions. This gives rise to a carbon-dependent consumer utility, which we define as $r_{D,d}-\lambda\cdot c_{D,d}$. Here, $r_{D,d}$ is the revenue for consumers derived from consuming electricity, $c_{D,d}$ denotes the carbon cost, given in units of [\$/tons $CO_{2}$] and $\lambda$ is the value of the average carbon signal. Specifically, given the price $p_{i}$ and the average carbon signal $\lambda$ calculated by ISO in the latter part, each consumer solves the following problem:
(2a) $\max_{P_{D,d}}\$$(r_{D,d}-p_{i:d\in\mathcal{D}_{i}}-\lambda\cdot c_{D,d})\cdot P_{D,d}$
(2b) $s.t.\$$P_{D,d}^{\min}\leq P_{D,d}\leq P_{D,d}^{\max},$
@@
where $p_{i:d\in\mathcal{D}_{i}}$ represents the electricity price at the node $i$ to which the consumer $d$ is connected. Each consumer incorporates the carbon term $\lambda\cdot c_{D,d}$ into their utility evaluation. In practice, the cost parameter $c_{D,d}$ may reflect concrete financial commitments, such as carbon taxes or the purchase of renewable energy certificates. It can also represent an internally defined “carbon cost”, reflecting a consumer’s willingness to forgo revenue to avoid carbon emissions. For example, if a consumer is willing to pay \$100/MWh for zero-carbon electricity but only \$50/MWh for electricity emitting 1 tCO2/MWh, the implied carbon cost is \$50/tCO2. Note that besides using carbon-dependent utilities to quantify the impacts of carbon emissions on consumers’ decision-making, there are also other ways to quantify this impact, e.g., setting carbon emission limits (Wang et al., 2021; Pietzcker et al., 2021). However, it can be very challenging for consumers to define what the values of the carbon emission cap should be.
@@
Transmission Owners: Transmission owners act as spatial arbitragers and aim to maximize the power transmission profit given electricity prices $p_{i}$. Specifically, the transmission owner maximizes their profit by buying power at a bus and selling it back at another one (arbitrage through the electricity price difference between two bus nodes), which is characterized by the following mathematical optimization problem:
(3a) $\max_{\theta_{i}}\$$\sum_{i\in\mathcal{N}}p_{i}\cdot\left(\sum_{j,(i,j)\in\mathcal{L}}\beta_{ij}(\theta_{j}-\theta_{i})\right)$
(3b) $s.t.\$$-F_{ij}^{\rm{lim}}\leq\beta_{ij}(\theta_{i}-\theta_{j})\leq F_{ij}^{\rm{lim}},\quad\forall(i,j)\in\mathcal{L},$
@@
where $\beta_{ij}\in\mathbb{R}$ denotes the susceptance value of the transmission line $(i,j)$ from node $i$ to node $j$. Constraints (3b) are the transmission line limits, where $F_{ij}^{\rm{lim}}$ represents the transmission capacity (which we assume is the same in both directions). Note that in the above, the voltage angle at the reference node $\theta_{\mathrm{ref}}$ is not a variable but the constant 0.
@@
ISO: The ISO clears the market by determining the electricity trading price at each node $i$ and provides the average carbon signal. Specifically, they set each nodal electricity price $p_{i}$ such that the supply-demand power balance
(4) $\quad\sum_{d\in\mathcal{D}_{i}}P_{D,d}+\sum_{j:(i,j)\in\mathcal{L}}\beta_{ij}(\theta_{i}-\theta_{j})=\sum_{g\in\mathcal{G}_{i}}P_{G,g}\ \qquad(\perp p_{i}),$
@@
(5) $\qquad\qquad\qquad\lambda\sum_{d\in\mathcal{D}}P_{D,d}=\sum_{g\in\mathcal{G}}e_{G,g}\cdot P_{G,g}\qquad\qquad(\perp\lambda),$
@@
where $e_{G}=\{e_{G,g}:\forall g\in\mathcal{G}\}$ is the carbon emission intensity vector for all generators (for generators using renewable sources, we assume that the carbon emission intensity values are zero).
@@
Note that (4) and (5) are written here as equality constraints but form a square complementarity system when paired with $p$ and $\lambda$, respectively. The number of equations in (4) (resp. (5)) matches the size of the variable $p$ (resp. $\lambda$). Both $p$ and $\lambda$ are parameters in the optimization problems (1)–(3).
The overall Nash equilibrium problem consists of the collection of parametric optimization problems (1), (2), (3) and the parametric complementarity conditions (4) and (5). We note that there exist alternative, equivalent formulations of this Nash equilibrium problem, which can be obtained by grouping optimality conditions differently. We show one such alternative formulation, which establishes a closer connection to the standard OPF market clearing, in Appendix A.
We assume that there exists an interior feasible point (no variable $P_{G,g},\forall g\in\mathcal{G}]$, $P_{D,d},\forall d\in\mathcal{D}]$, and $\theta_{i},\forall i\in\mathcal{N}\setminus{\rm ref}]$ are at their bounds) such that the constraints  (4) and $\theta_{\rm ref}=0$ are satisfied. This ensures the existence of a solution to our equilibrium problem, see Appendix C.
@@
When all $c_{D}=0$, our equilibrium problem reduces to the standard social welfare maximization problem. The equilibrium framework seeks to coordinate all participants toward a consistent set of values for $P_{G}$, $P_{D}$, $p$, and $\lambda$ that simultaneously ensures power balance and aligns with the optimal solutions of each participant’s individual optimization problem.
Unlike the sequential method, where inaccuracies may arise from relying on lagged market-clearing signals (see (Gorka et al., 2025; Lindberg et al., 2022, 2021a)), the equilibrium method uses synchronized decision-making based on perfect real-time carbon signals (as demonstrated by the alternative formulation in Appendix A), which we believe is better for analyzing the effectiveness of different carbon signals. We note that, similar to the standard (carbon-agnostic) market-clearing optimization problem, the equilibrium problem may have multiple solutions (e.g., when multiple generators at a node share identical costs).
2.2. Three-bus Example Illustration